x + 2y = 14
y = 3x - 14

Answer by Techpriest(29)

Your problem that you have gave out may be solved through the method of substitution.
Plug in the y from the second equation to the first equation.
x + 2(3x-14) = 14
x + 6x - 28 = 14 Distribution
7x - 28 = 14 Combine like terms
7x = 42 Addition Property of Equality
7x/7 = 42/7 Division Property of Equality
x = 6 Simplify
When we check our answers...
6 + 2(6(3)-14) = 14 Substitution
6 + 2(18-14) = 14 Simplify
6 + 2(4) = 14 Simplify
6 + 8 = 14 Simplify
14 = 14 Simplify